# Correlations versus noise in the NFT market

### v1

@@ Abstract
The non-fungible token (NFT) market emerges as a recent trading innovation leveraging blockchain technology, mirroring the dynamics of the cryptocurrency market. To deepen the understanding of the dynamics of this market, in the current study, based on the capitalization changes and transaction volumes across a large number of token collections on the Ethereum platform, the degree of correlation in this market is examined by using the multivariate formalism of detrended correlation coefficient and correlation matrix. It appears that correlation strength is lower here than that observed in previously studied markets. Consequently, the eigenvalue spectra of the correlation matrix more closely follow the Marchenko-Pastur distribution, still, some departures indicating the existence of correlations remain. The comparison of results obtained from the correlation matrix built from the Pearson coefficients and, independently, from the detrended cross-correlation coefficients suggests that the global correlations in the NFT market arise from higher frequency fluctuations. Corresponding minimal spanning trees (MSTs) for capitalization variability exhibit a scale-free character while, for the number of transactions, they are somewhat more decentralized.
The complexity of financial markets, encompassing traditional assets and emerging forms like cryptocurrencies and non-fungible tokens (NFTs), is a multi-scale phenomenon driven by various factors. Understanding this complexity and the associated emergent phenomena requires delving into the details and intricacies of these markets, dynamic interaction of diverse elements, and the evolving system of global finance. NFTs represent a unique form of digital assets, often tied to digital art, collectibles, or virtual real estate. The subtleties related to complexity of NFT markets arise from their innovative nature resulting from the valuation of digital assets, intellectual property considerations, and the variety of smart contracts. NFTs also introduce new possibilities for creators and collectors, but come with challenges related to copyright, authenticity, and market saturation. The issue of global correlations within NFT markets has not been quantitatively addressed in the literature so far and therefore, in this study, it is framed in the context of the potential competition between correlation effects and randomness.
@@ I Introduction
The cryptoassets space is constantly evolving. Since the first cryptocurrency - Bitcoin - launched in 2009 Nakamoto (2009), blockchain technology has undergone considerable evolution and found applications in various fields Sunny et al. (2022). The initial functionality of Bitcoin and its early alternatives was limited to settling transactions and serving as a distributed ledger register. This changed with the launch of Ethereum in 2015 as a decentralized, open-source, and distributed computing platform ETH . This marked the first significant expansion of blockchain capabilities. The Ethereum smart contracts functionality allowed for the easy creation of one’s own fungible tokens in ERC-20 standard ERC-20 protocol . One of the applications of smart contracts was one of the first and highest-valued NFT collections - CryptoPunks, even though the concept of unique tokens is older and can be traced back to 2012 Bitcoin colored coins Rosenfeld (2012) and the creation of the first-ever NFT, Quantum, minted by Kevin McCoy on Namecoin in 2014 quantmNFT . The Ethereum project experienced significant growth and development, particularly in 2017, when numerous projects launched on its blockchain through initial coin offerings (ICOs), ultimately contributing to a market bubble Bellón and Figuerola-Ferretti (2022). At the peak of the ICO bubble in December 2017, the challenges in the capacity of the Ethereum network became increasingly evident. The surge in popularity of the CryptoKitties game exacerbated these issues, at times consuming up to 70% of Ethereum’s usage capacity. This spike in demand led to record highs in the number of transactions and fees, considerably slowing down the network Jiang and Liu (2021). The game enables participants to purchase, sell, and create non-fungible tokens (NFTs) on the Ethereum blockchain. It stands out as one of the earliest and most prominent examples of the ERC-721 standard, which became officially recognized as the Non-Fungible Token standard in 2018 ERC-721 protocol . This standard facilitates the straightforward creation of NFT tokens and collections by anyone.
@@
The application of smart contracts to finance marked a significant second milestone for the cryptoasset domain, giving rise to the phenomenon known as DeFi (Decentralized Finance) Summer that lasted between mid-2020 and early 2021 Maouchi, Charfeddine, and El Montasser (2022). This was accompanied by the development of blockchains that competed with Ethereum, such as Solana, Avalanche, Arbitrum, BNB Chain, Polygon, Tron, Optimism, and Cardano. The total value locked in DeFi protocols on various blockchains peaked in Dec 2021 with a value of 250 billion USD TVL . In this period, NFT tokens surged in popularity and entered the mainstream, driven by a low barrier to entry, the simplicity of launching personal collections, and celebrity endorsements. These factors fueled the speculative NFT bubble of 2021 Mingjun Guo and Wei (2023). The boom started in March 2021 when Mike Winkelmann, aka Beeple, sold his NFT Everydays: The First 5,000 Days (2021) for \$69 million at Christie’s Christies Press Center . Shortly after that, the 10,000-part series Bored Ape Yacht Club was launched in May 2021, starting a wave of similar projects called profile-pic NFTs (PFP NFTs). Many celebrities owed a Bored Ape Yacht Club NFT or released their own collections Horky et al. (2023); Kim and Chakraborty (2024); NFTceleb and even the former US president Donald Trump had participated in the NFT collective enthusiasm TrumpNFT . After spectacular growth, when the most expensive NFT token (CryptoPunk #$\#5822$) was sold for 8000 ETH in February 2022, the NFT market eventually peaked in May 2022, reaching a total capitalization of $\sim$21 billion USD NFTbubble . The widespread popularity of NFTs at that time allowed for funding support for Ukraine via selling one of the CryptoPunk tokens in June 2022 for 90 ETH ($\sim$100,000 USD). After the subsequent crash NFTcrash, , the total capitalization dropped to 5 billion USD in April 2024 coinmarketcap . Given the unregulated nature of the NFT market, it has been subject to numerous manipulative practices Bose et al. (2023); Sifat, Tariq, and van Donselaar (2024) including wash trading von Wachter et al. (2022) and pump-and-dump schemes Bose et al. (2024).
@@
Despite a dramatic drop of nearly 90% in market value, the blockchain technology underpinning NFTs continues to evolve. The Bitcoin Ordinals protocol was introduced to the Bitcoin network in January 2023 BTCordinals , enabling the creation of collections on the Bitcoin blockchain. In December 2023, this blockchain accounted for the largest portion of NFT trading volume at 42.1%. Nonetheless, throughout 2023, Ethereum maintained its position as the dominant blockchain for NFT transactions with 72.3% trading volume share Blokchianshare . Other blockchains hosting NFT collections include Solana, Immutable X, Polygon, BNB Chain, Flow, Arbitrum, Avalanche, and Ronin. There are many NFT marketplaces that enable NFT collections trading; the three largest ones are Blur, Magic Eaden and OpenSea as of April 2024 NFTmarketplaces . Although the current main NFT application to pure speculation faced serious criticism NFTcrit and, despite many challenges Ali et al. (2023), the standard may still represent a paradigm shift in how digital assets are perceived, owned, and transacted, reshaping the digital world’s understanding of value and ownership Chalmers et al. (2022); Giannoni, Medda, and Bartolucci (2023).
@@
Building on our previous exploration of price fluctuation characteristics within individual NFT collections Szydło et al. (2024), this article aims to delve into the correlations among NFT collection trades. The well-established financial markets, such as stocks James and Menzies (2021); James, Menzies, and Gottwald (2022), Forex Drożdż, Górski, and Kwapień (2007); Gębarowski et al. (2019); Miśkiewicz (2021), and even younger cryptocurrencies Stosic et al. (2018); Chaudhari and Crane (2020); Drożdż et al. (2020); Wątorek et al. (2021); James and Menzies (2022); Nguyen et al. (2022); Gavin and Crane (2023); James and Menzies (2023), demonstrate significant collective behavior in their cross-correlations. Random Matrix Theory (RMT) has been effectively utilized to uncover the nature and origins of correlations in these markets Laloux et al. (1999); Plerou et al. (1999); Drożdż et al. (2001). In this study, RMT is applied to examine the price and the number of transactions variations of the collections created on the Ethereum blockchain. This analysis will focus on determining the non-random nature of these correlations and examining if distinct NFT-market sectors can be distinguished. Properties of minimal spanning trees constructed from the correlation networks based on correlation matrices will also be analyzed. This approach has already been successfully applied to financial time series Mantegna (1999); Tumminello, Lillo, and Mantegna (2010), including stocks Miccichè et al. (2003); Wiliński et al. (2013); Miśkiewicz and Bonarska-Kujawa (2022), currencies Górski, Drożdż, and Kwapień (2008); Gębarowski et al. (2019), and cryptocurrencies Polovnikov, Kazakov, and Syntulsky (2020); Wątorek et al. (2021); Drożdż, Kwapień, and Wątorek (2023); Brigatti et al. (2023); Nguyen et al. (2023).
@@ II.1 Data specification
Understanding the characteristics of time series is crucial before proceeding with a correlation analysis, because this knowledge can help one in choosing the right correlation metrics. For this purpose, cumulative distributions of the capitalization increments $c_{\Delta t}$ and the number of transactions $N_{\Delta t}$ were analyzed for the 13 most liquid collections (those with %$\%0_{\Delta t=1{\rm h}}<50\%$ in Table 1). Fig. 3 shows that the cumulative probability distribution functions (CDFs) $P(X>\sigma)$, where $\sigma$ denotes standard deviation, of $|c_{\Delta t}|$ and $N_{\Delta t}$ for all the collections exhibit heavy tails – a common characteristics of financial time series Ausloos (2000); Cont (2001); Plerou et al. (2004); Wątorek, Kwapień, and Drożdż (2021).
For the absolute capitalization increments $|c_{\Delta t}|$, the CDF tails follow a power law $x^{-\gamma}$ with the exponent $\gamma$ varying between 1.5 and 2.0 for most collections – see Fig. 3(a). For the most liquid ones, such as Bored Ape Yacht Club and Mutant Ape Yacht Club, a stretched exponential function ${\rm exp}(x^{\beta})$ provides a good fit if $\beta\approx 0.35$. In terms of the number of transactions $N_{\Delta t}$, the CDF tails are less heavy as compared to those for $|c_{\Delta t}|$, with the stretched exponential function accurately fitting the data for $0.3\leq\beta\leq 0.4$. These findings align closely with the results from a previous study on the most liquid collections from the Solana blockchain Szydło et al. (2024), where similar patterns were observed in the distributions of absolute capitalization increments and the number of transactions.
Figure 3: Cumulative probability distribution functions for (a) absolute logarithmic increments of collection capitalization expressed in USD $|c_{\Delta t=1h}(t)|$ and (b) the number of transactions aggregated hourly $N_{\Delta t=1h}(t)$.
Another measure of temporal relationships in time series, which may involve both linear and non-linear dependencies, is the autocorrelation function
@@ II.2 Detrended correlations
where the polynomials $P^{(m)}$ of order $m$ are utilized. Specifically, an order of $m=2$ was chosen for its effectiveness with financial time series Oświęcimka et al. (2013). Following this procedure, a total of $2M_{s}$ segments containing detrended signals are generated. Subsequently, the variance and covariance for each segment $\nu$ are calculated as follows:
@@
Figure 5: Fluctuation functions $F(q,s)$ with $q\in[-4,4]$ calculated for the number of transactions $N_{\Delta t=1h}$ and logarithmic increments of the total collection capitalization $c_{\Delta t=1h}$ for Mutant Ape Yacht Club collection. In main panels, thick green lines represent $F(q=2,s)$, which are utilized in computation of the Hurst exponent $H$. Vertical red dashed lines point out to a scale range where the family of $F(q,s)$ exhibit power-law dependence for different values of $q$, from which the singularity spectrum $f(\alpha)$ can be calculated (insets).
@@ III.1 Eigenvalues and off-diagonal elements distributions
The market factor becomes weaker, although still significant, if the matrices ${\bf C}^{\rho}(q,s)$ are considered (middle panels of Fig 6(a) for $q=1$ and bottom panels for $q=4$). In this case, one could also observe an effect of the correlations increasing with time scale $s$, which is typical for the financial markets Wątorek et al. (2019). The probability distribution function of the off-diagonal elements of the correlation matrices ${\bf C}$ and ${\bf C}^{\rho}(q,s)$ offers complementary information to the eigenvalue distribution. Deviations from a Gaussian distribution in this case are visible in all the cases in Fig 6(b). The fattest tails are observed for $q=4$ (large fluctuations) for $N_{\Delta t}$ see bottom left panel of Fig 6(b), where the extreme elements $C_{ij}^{\rho}(q,s)$ exceed 0.9 even on the shortest time scale $s=7$ days. These large correlations can be traced back to a number of transactions surge across various less-traded collections on the same day Jan, 6 2023 (inset of Fig. 1(b)). Their origin will be explained in the following section.
Figure 6: (a) Eigenvalue distributions of the correlation matrices ${\bf C}$ (top) and ${\bf C}^{\rho}(q,s)$ (middle for $q=1$ and bottom for $q=4$) created from the total collection capitalization increments $c_{\Delta t}$ (left) and the number of transactions $N_{\Delta t}$ (right) for $\Delta t=1$h. The theoretical Marchenko–Pastur distribution $\phi_{\rm W}(\lambda)$ is marked by dashed red lines. (b) Probability density functions of the off-diagonal elements of the same correlation matrices and data as in (a). Dashed lines show a fitted normal distribution to each empirical distribution (histograms).

### v2

@@ Abstract
The non-fungible token (NFT) market emerges as a recent trading innovation leveraging blockchain technology, mirroring the dynamics of the cryptocurrency market. The current study is based on the capitalization changes and transaction volumes across a large number of token collections on the Ethereum platform. In order to deepen the understanding of the market dynamics, the collection-collection dependencies are examined by using the multivariate formalism of detrended correlation coefficient and correlation matrix. It appears that correlation strength is lower here than that observed in previously studied markets. Consequently, the eigenvalue spectra of the correlation matrix more closely follow the Marchenko-Pastur distribution, still, some departures indicating the existence of correlations remain. The comparison of results obtained from the correlation matrix built from the Pearson coefficients and, independently, from the detrended cross-correlation coefficients suggests that the global correlations in the NFT market arise from higher frequency fluctuations. Corresponding minimal spanning trees (MSTs) for capitalization variability exhibit a scale-free character while, for the number of transactions, they are somewhat more decentralized.
The complexity of financial markets, encompassing traditional assets and emerging forms like cryptocurrencies and non-fungible tokens (NFTs), is a multi-scale phenomenon driven by various factors. Understanding this complexity and the associated emergent phenomena requires delving into the details and intricacies of these markets, dynamic interaction of diverse elements, and the evolving system of global finance. NFTs represent a unique form of digital assets, which may be tied to digital art, collectibles, or virtual real estate. The additional complexity of the NFT markets arise from the innovative aspects of digital asset valuation, intellectual property considerations, and the other aspects of smart contracts. NFTs also introduce new possibilities for creators and collectors, but come with challenges related to copyright, authenticity, and market saturation. The issue of global correlations within NFT markets has not been quantitatively addressed in the literature so far and therefore, in this study, it is framed in the context of the potential competition between correlation effects and randomness.
@@ I Introduction
The cryptoassets space is constantly evolving. Since the first cryptocurrency - Bitcoin - launched in 2009 Nakamoto (2009), blockchain technology has undergone considerable evolution and found applications in various fields Sunny et al. (2022). The initial functionality of Bitcoin and its early alternatives was limited to settling transactions and serving as a distributed ledger register. This changed with the launch of Ethereum in 2015 as a decentralized, open-source, and distributed computing platform ETH . This marked the first significant expansion of blockchain capabilities. The Ethereum smart contracts functionality allowed for the easy creation of one’s own fungible tokens in ERC-20 standard ERC-20 protocol . Among the early applications of smart contracts was one of the first and highest-valued NFT collections - CryptoPunks. However, the concept of unique tokens is older and can be traced back to Bitcoin colored coins in 2012 Rosenfeld (2012) and the creation of the first-ever NFT, Quantum, minted by Kevin McCoy on Namecoin in 2014 quantmNFT . The Ethereum project experienced significant growth and development, particularly in 2017, when numerous projects launched on its blockchain through initial coin offerings (ICOs), ultimately contributing to a market bubble Bellón and Figuerola-Ferretti (2022). At the peak of the ICO bubble in December 2017, the challenges in the capacity of the Ethereum network became increasingly evident. The surge in popularity of the CryptoKitties game exacerbated these issues, at times consuming up to 70% of Ethereum’s usage capacity. This spike in demand led to record highs in the number of transactions and fees, considerably slowing down the network Jiang and Liu (2021). The game enables participants to purchase, sell, and create non-fungible tokens (NFTs) on the Ethereum blockchain. It stands out as one of the earliest and most prominent examples of the ERC-721 standard, which became officially recognized as the Non-Fungible Token standard in 2018 ERC-721 protocol . This standard facilitates the straightforward creation of NFT tokens and collections by anyone.
@@
The application of smart contracts to finance marked a significant second milestone for the cryptoasset domain, giving rise to the phenomenon known as DeFi (Decentralized Finance) Summer that lasted between mid-2020 and early 2021 Maouchi, Charfeddine, and El Montasser (2022). This was accompanied by the development of blockchains that competed with Ethereum, such as Solana, Avalanche, Arbitrum, BNB Chain, Polygon, Tron, Optimism, and Cardano. The total value locked in DeFi protocols on various blockchains peaked in Dec 2021 with a value of 250 billion USD TVL . In this period, NFT tokens surged in popularity and entered the mainstream, driven by a low barrier to entry, the simplicity of launching personal collections, and celebrity endorsements. These factors fueled the speculative NFT bubble of 2021 Mingjun Guo and Wei (2023). The boom started in March 2021 when Mike Winkelmann, aka Beeple, sold his NFT Everydays: The First 5,000 Days (2021) for \$69 million at Christie’s Christies Press Center . Shortly after that, the 10,000-part series Bored Ape Yacht Club was launched in May 2021, starting a wave of similar projects called profile-pic NFTs (PFP NFTs). Many celebrities owned a Bored Ape Yacht Club NFT or released their own collections Horky et al. (2023); Kim and Chakraborty (2024); NFTceleb and even the former US president Donald Trump had participated in the NFT collective enthusiasm TrumpNFT . After spectacular growth, when the most expensive NFT token (CryptoPunk #$\#5822$) was sold for 8000 ETH in February 2022, the NFT market eventually peaked in May 2022, reaching a total capitalization of $\sim$21 billion USD NFTbubble . The widespread popularity of NFTs at that time allowed for funding support for Ukraine via selling one of the CryptoPunk tokens in June 2022 for 90 ETH ($\sim$100,000 USD). After the subsequent crash NFTcrash, , the total capitalization dropped to 5 billion USD in April 2024 coinmarketcap . Given the unregulated nature of the NFT market, it has been subject to numerous manipulative practices Bose et al. (2023); Sifat, Tariq, and van Donselaar (2024) including wash trading von Wachter et al. (2022) and pump-and-dump schemes Bose et al. (2024).
@@
Despite a dramatic drop of nearly 90% in market value, the blockchain technology underpinning NFTs continues to evolve. The Bitcoin Ordinals protocol was introduced to the Bitcoin network in January 2023 BTCordinals , enabling the creation of collections on the Bitcoin blockchain. In December 2023, this blockchain accounted for the largest portion of NFT trading volume at 42.1%. Nonetheless, throughout 2023, Ethereum maintained its position as the dominant blockchain for NFT transactions with 72.3% trading volume share Blokchianshare . Other blockchains hosting NFT collections include Solana, Immutable X, Polygon, BNB Chain, Flow, Arbitrum, Avalanche, and Ronin. There are many NFT marketplaces that enable NFT collections trading; the three largest ones are Blur, Magic Eaden and OpenSea as of April 2024 NFTmarketplaces . Although the current main NFT application to pure speculation faced serious criticism NFTcrit and, despite many challenges Ali et al. (2023), the standard may still bring a new outlook in how digital assets are perceived, owned, and transacted, reshaping the digital world’s understanding of value and ownership Chalmers et al. (2022); Giannoni, Medda, and Bartolucci (2023).
@@
Building on our previous exploration of price fluctuation characteristics within individual NFT collections Szydło et al. (2024), this article aims to delve into the correlations among NFT collection trades. The well-established financial markets, such as stocks James and Menzies (2021); James, Menzies, and Gottwald (2022), Forex Drożdż, Górski, and Kwapień (2007); Gębarowski et al. (2019); Miśkiewicz (2021), and even younger cryptocurrencies Stosic et al. (2018); Chaudhari and Crane (2020); Drożdż et al. (2020); Wątorek et al. (2021); James and Menzies (2022); Nguyen et al. (2022); Gavin and Crane (2023); James and Menzies (2023), demonstrate significant collective behavior in their cross-correlations. Random Matrix Theory (RMT) has been effectively utilized to uncover the nature and origins of correlations in these markets Laloux et al. (1999); Plerou et al. (1999); Drożdż et al. (2001). In this study, RMT is applied to examine the price and the number of transactions variations of the collections created on the Ethereum blockchain. This analysis will focus on determining the non-random nature of these correlations and examining if distinct NFT-market sectors can be distinguished. Properties of minimal spanning trees constructed from the correlation networks will also be analyzed. This approach has already been successfully applied to financial time series Mantegna (1999); Tumminello, Lillo, and Mantegna (2010), including stocks Miccichè et al. (2003); Wiliński et al. (2013); Miśkiewicz and Bonarska-Kujawa (2022), currencies Górski, Drożdż, and Kwapień (2008); Gębarowski et al. (2019), and cryptocurrencies Polovnikov, Kazakov, and Syntulsky (2020); Wątorek et al. (2021); Drożdż, Kwapień, and Wątorek (2023); Brigatti et al. (2023); Nguyen et al. (2023).
@@ II.1 Data specification
Understanding the characteristics of time series is crucial before proceeding with a correlation analysis, because this knowledge can help one in choosing the right correlation metrics. For this purpose, cumulative distributions of the capitalization increments $c_{\Delta t}$ and the number of transactions $N_{\Delta t}$ were analyzed for the 13 most liquid collections (those with %$\%0_{\Delta t=1{\rm h}}<50\%$ in Table 1). Fig. 3 shows that the complementary cumulative distribution functions (CCDFs) $P(X>\sigma)$, where $\sigma$ denotes the estimate of standard deviation, of $|c_{\Delta t}|$ and $N_{\Delta t}$ for all the collections exhibit heavy tails – a common characteristics of financial time series Ausloos (2000); Cont (2001); Plerou et al. (2004); Wątorek, Kwapień, and Drożdż (2021).
For the absolute capitalization increments $|c_{\Delta t}|$, the CCDF tails follow a power law $x^{-\gamma}$ with the exponent $\gamma$ varying between 1.5 and 2.0 for most collections – see Fig. 3(a). For the most liquid ones, such as Bored Ape Yacht Club and Mutant Ape Yacht Club, a stretched exponential function ${\rm exp}(x^{\beta})$ provides a good fit if $\beta\approx 0.35$. In terms of the number of transactions $N_{\Delta t}$, the CCDF tails are less heavy as compared to those for $|c_{\Delta t}|$, with the stretched exponential function accurately fitting the data for $0.3\leq\beta\leq 0.4$. These findings align closely with the results from a previous study on the most liquid collections from the Solana blockchain Szydło et al. (2024), where similar patterns were observed in the distributions of absolute capitalization increments and the number of transactions.
Figure 3: Complementary cumulative distribution functions for (a) absolute logarithmic increments of collection capitalization expressed in USD $|c_{\Delta t=1h}(t)|$ and (b) the number of transactions aggregated hourly $N_{\Delta t=1h}(t)$.
Another measure of temporal relationships in time series is the autocorrelation function
@@ II.2 Detrended correlations
by using ordinary least-squares method and the polynomials $P^{(m)}$ of order $m$ are utilized. Specifically, an order of $m=2$ was chosen for its effectiveness with financial time series Oświęcimka et al. (2013). Following this procedure, a total of $2M_{s}$ segments containing detrended signals are generated. Subsequently, the estimate of variance and estimate of covariance for each segment $\nu$ are calculated as follows:
@@
Figure 5: Fluctuation functions $F(q,s)$ with $q\in[-4,4]$ calculated for the number of transactions $N_{\Delta t=1h}$ and logarithmic increments of the total collection capitalization $c_{\Delta t=1h}$ for Mutant Ape Yacht Club collection. In main panels, thick green lines represent $F(q=2,s)$, which are utilized in the estimation of the Hurst exponent $H$ with the standard error. Vertical red dashed lines point out to a scale range where the family of $F(q,s)$ exhibit power-law dependence for different values of $q$, from which the singularity spectrum $f(\alpha)$ can be calculated (insets).
@@ III.1 Eigenvalues and off-diagonal elements distributions
The market factor becomes weaker, although still significant, if the matrices ${\bf C}^{\rho}(q=1,s)$ and ${\bf C}^{\rho}(q=4,s)$ are considered (middle and bottom panels of Fig 6(a)). In this case, one could also observe an effect of the correlations increasing with time scale $s$, which is typical for the financial markets Wątorek et al. (2019). The probability distribution function (histograms) of the off-diagonal elements of the correlation matrices ${\bf C}$ and ${\bf C}^{\rho}(q,s)$ offers complementary information to the eigenvalue distribution. Deviations from a Gaussian distribution in this case are visible in all the cases in Fig 6(b). The fattest tails are observed for $q=4$ (large fluctuations) for $N_{\Delta t}$ see bottom left panel of Fig 6(b), where the extreme elements $C_{ij}^{\rho}(q=4,s)$ exceed 0.9 even on the shortest time scale $s=7$ days. These large correlations can be traced back to a number of transactions surge across various less-traded collections on the same day Jan, 6 2023 (inset of Fig. 1(b)). Their origin will be explained in the following section.
Figure 6: (a) Eigenvalue distributions of the correlation matrices ${\bf C}$, ${\bf C}^{\rho}(q=1,s)$ and ${\bf C}^{\rho}(q=4,s)$ (top, middle, bottom) created from the total collection capitalization increments $c_{\Delta t}$ (left) and the number of transactions $N_{\Delta t}$ (right) for $\Delta t=1$h. The theoretical Marchenko–Pastur distribution $\phi_{\rm W}(\lambda)$ is marked by dashed red lines. (b) Probability distribution functions (histograms) of the off-diagonal elements of the same correlation matrices and data as in (a). Dashed lines show a fitted normal distribution to each empirical distribution.